### v2

@@ 2.2 Analyzing the linear program for the Lovász number of generalized Johnson graphs
$\sum_{\lambda\in\Lambda}\text{sgn}(\lambda)\prod_{i=m+1}^{m+p}P_{s_{(i,\lambda(i))}}=(-1)^{p}\frac{\prod_{i=1}^{p}(\ell_{m+i}+1)}{p!\prod_{i=1}^{p}(k-\ell_{m+i}-1)!}n^{pk-(\sum_{i=1}^{p}\ell_{m+i})-p}+O(n^{pk-(\sum_{i=1}^{p}\ell_{m+i})-p-1}).$
@@
We proceed by strong induction on $p$. The case $p=1$ follows from (5), as $\lambda(m+1)=m+1$, so $\text{sgn}(\lambda)=1$ and $P_{s_{(m+1,\lambda(m+1))}}=(-1)\frac{\ell_{m+1}}{k-\ell_{m+1}-1}n^{k-\ell_{m+1}-1}+O(n^{k-\ell_{m+1}-2})$. Now suppose the statement of the lemma holds for all $11\leq r\leq p-$ 1; that is, assume that the lemma holds for any run of $r$ consecutive integers $\{\ell_{y},\ell_{y+1},\ldots,\ell_{y+r-1}\}$ in $L$. We show that the statement holds for $L_{m}$ under this assumption. Divide the permutations $\lambda$ in $\Lambda$ up according to the value of $j$ for which $\lambda(m+j)=m+1$. It follows from the conditions on $\lambda$ that $\lambda(m+i)=m+i+1$ for $1\leq i<j$, so that for these permutations $\lambda$,
$\prod_{i=m+1}^{m+j}P_{s_{(i,\lambda(i))}}=\binom{\ell_{m+1}+j}{\ell_{m+1}}(-1)^{\ell_{m+j}+1-\ell_{m+1}}\prod_{i=1}^{j}\frac{1}{(k-\ell_{m+i}-1)!}n^{jk-\sum_{i=1}^{j}\ell{m+i}-j}+O(n^{jk-\sum_{i=1}^{j}\ell{m+i}-j})$
$=(-1)^{j}\frac{\prod_{i=1}^{j}(\ell_{m+i}+1)}{j!\prod_{i=1}^{j}(k-\ell_{m+i}-1)!}n^{jk-\sum_{i=1}^{j}\ell{m+i}-j}+O(n^{jk-\sum_{i=1}^{j}\ell{m+i}-j}).$
@@
$\sum_{\begin{subarray}{c}\lambda\in\Lambda\\ \lambda(m+j)=m+1\end{subarray}}\text{sgn}(\lambda_{[m+j+1,m+p]})\prod_{i=m+j+1}^{m+p}P_{s_{(i,\lambda(i))}}$
$=(-1)^{p-j}\frac{1}{(p-j)!}\frac{\prod_{i=j+1}^{p}(\ell_{m+i}+1)}{\prod_{i=j+1}^{p}(k-\ell_{m+i}-1)!}n^{(p-j)k-(\sum_{i=j+1}^{p}\ell_{m+i})-(p-j)-1}+O(n^{(p-j)k-(\sum_{i=j+1}^{p}\ell_{m+i})-(p-j)-1}),$
@@
$\sum_{\begin{subarray}{c}\lambda\in\Lambda\\ \lambda(m+j)=m+1\end{subarray}}\text{sgn}(\lambda)\prod_{i=m+j}^{m+p}P_{s_{(i,\lambda(i))}}$
@@
$(-1)^{p-j-1}\frac{1}{j!(p-j)!}\frac{\prod_{i=1}^{p}(\ell_{m+i}+1)}{\prod_{i=1}^{p}(k-\ell_{m+i}-1)!}n^{pk-(\ell_{m+1}+\cdots+\ell_{m+p})-p}+O(n^{pk-(\ell_{m+1}+\cdots+\ell_{m+p})-p-1}).$
@@
$\sum_{\lambda\in\Lambda}\text{sgn}(\lambda)\prod_{i=m+1}^{m+p}P_{s_{(i,\lambda(i))}}=\sum_{j=1}^{p}\sum_{\begin{subarray}{c}\lambda\in\Lambda\\ \lambda(m+j)=m+1\end{subarray}}\text{sgn}(\lambda)\prod_{i=m+1}^{m+p}P_{s_{(i,\lambda(i))}}$
@@
$\sum_{\sigma\in\Sigma}\text{sgn}(\sigma)\prod_{i=1}^{s}P_{s_{(i,\sigma(i))}}=\prod_{j=1}^{b}\left(\sum_{\sigma_{r_{i}}}\text{sgn}(\sigma_{r_{i}})\prod_{a=1}^{m_{i}}P_{s_{(a,\sigma_{r_{i}}(a))}}\right)$
@@
For each $1\leq i\leq s$, let $a_{k-\ell_{i}}$ be the expression given by (4). Then, $a_{k-\ell_{i}}$ is a polynomial in $n$ of degree $s-i+1$ with positive leading order term $C_{i}$, so that
@@
${\bf v}=P_{s}^{-1}\begin{pmatrix}0\\ 0\\ \vdots\\ 0\\ -1\end{pmatrix}.$
Therefore, we only need to analyze the terms in column $s$ of $P_{s}^{-1}$. We first determine the leading order term of $P_{s_{(1,s)}}^{-1}$. We have $P_{s_{(1,s)}}^{-1}=\frac{(-1)^{s+1}}{\text{det}(P_{s})}\text{det}(P_{s}^{(s,1)})$. We find the leading order term of $\text{det}(P_{s}^{(s,1)})$ by the Leibniz formula. By (5), the leading order term in the first row is $\frac{1}{(k-\ell_{2})!}\binom{k-\ell_{1}-1}{k-\ell_{2}}n^{k-\ell_{2}}$, as $P_{s_{(1,i)}}^{(s,1)}=\Theta(n^{k-\ell_{i+1}})$ for $11\leq i\leq s-$ 1. Inductively, the leading order term in the $j$th row is $\frac{\binom{k-\ell_{j-1}-1}{k-\ell_{j}}}{(k-\ell_{j})!}n^{k-\ell_{j}}$, so
$\text{det}(P_{s}^{(s,1)})=\binom{k-\ell_{1}-1}{k-\ell_{2}}\cdots\binom{k-\ell_{s-1}-1}{k-\ell_{s}}\frac{1}{\prod_{i=2}^{s}(k-\ell_{i})!}n^{(s-1)k-\sum_{i=2}^{s}\ell_{i}}+O(n^{(s-1)k-\sum_{i=2}^{s}\ell_{i}-1}).$
@@
$a_{k-\ell_{1}}=\frac{(-1)\binom{k}{k-\ell_{1}}\binom{n-k}{k-\ell_{1}}(-1)^{s+1}\text{det}(P_{s}^{(s,1)})}{\text{det}(P_{s})}$
@@
Now, let $2\leq i\leq s$. We find the leading order term of $\text{det}(P_{s}^{(s,i)})$ by the Leibniz formula. Consider the term in row $i-1$ from any permutation $\sigma$ which contributes a leading order term to the Leibniz formula. If $\sigma(i-1)\geq i$ (in $P_{s}^{(s,i)}$), then by (5), the term contributed by $\sigma$ to the Leibniz formula will be of smaller order. Hence, restricted to the set $[1,i-1]$, the permutations $\sigma$ which contribute to the leading order term of $\text{det}(P_{s}^{(s,i)})$ permute the set $[1,i-1]$ according to the restrictions of Claim 2.9. Hence, by the proof of Lemma 2.8, these terms contribute $(-1)^{i-1}D_{i}n^{k(i-1)-(\ell_{1}+\cdots+\ell_{i-1})-(i-1)}+O(n^{k(i-1)-(\ell_{1}+\cdots+\ell_{i-1})-(i-1)-1})$ to $\text{det}(P_{s}^{(s,i)})$, where $D_{i}$ is a positive constant whose exact value we omit. For rows $i$ through $s-1$ in $P_{s}^{(s,i)}$, the leading order terms are $P_{s_{(m,m)}}^{(s,i)}$ for $i\leq m\leq s-1$, which contribute $E_{i}n^{(s-i)k-(\ell_{i+1}+\ldots+\ell_{s})}+O(n^{(s-i)k-(\ell_{i+1}+\ldots+\ell_{s})-1})$, where $E_{i}$ is a positive constant whose exact value we omit. These calculations give that
$a_{k-\ell_{i}}=\frac{(-1)\binom{k}{k-\ell_{i}}\binom{n-k}{k-\ell_{i}}}{\text{det}(P_{s})}\left((-1)^{s+i}(-1)^{i-1}D_{i}E_{i}n^{(s-1)k-(\sum_{j=1,j\neq i}^{s}\ell_{j})-(i-1)}+O(n^{(s-1)k-(\sum_{j=1,j\neq i}^{s}\ell_{j})-(i-1)-1})\right)$
$=C_{i}n^{s-i+1}-O(n^{s-i}),$
where $C_{i}$ is a positive constant. (The constants $C_{i}$ can be determined from the calculations presented here, but we omit them because they are of lower order). ∎
@@
The inequalities (8) with $\ell_{m}+2\leq u\leq\ell_{m+1}$ for $11\leq m\leq s-$ 1 will be satisfied for large $n$ as the leading order term on the left-hand side is $a_{k-\ell_{m+1}}\frac{\binom{k-u}{k-\ell_{m}}}{\binom{k}{k-\ell_{m}}}=C_{m+1}\frac{\binom{k-u}{k-\ell_{m}}}{\binom{k}{k-\ell_{m}}}n^{s-m}$.
@@
It would be interesting to strengthen Theorem 1.2 by finding the exact value of $\vartheta(G(n,k,L))$ (as Theorem 1.5 does if $|L|=1$ or $|L|=k-1$) as $n\rightarrow\infty$. We have opted for the computations presented here, as they were sufficient to obtain the order of magnitude of $\vartheta(G(n,k,L))$ and the correct constant in front of the $n^{s}$ term.
We note that the argument used to prove Theorem 1.2 also asymptotically determines the best possible result the Delsarte linear programming bound [Del73, Sch79] can give for the graphs $G(n,k,L)$. Since the graphs $G(n,k,L)$ are unions of graphs in the Johnson scheme, Delsarte’s linear programming bound $\sigma(G(n,k,L))$ is the same linear program as the linear program for the Lovász number in (1), with the additional restriction that the variables $a_{0},a_{1},\ldots,a_{n}$ must all be nonnegative. The solution in the proof of Theorem 1.2 has all $a_{k-\ell_{i}}$ nonnegative, so (4) is also a feasible solution for Delsarte’s linear programming bound, so as $n\rightarrow\infty$,
@@
Let $G=G(n,k,L)$ be a generalized Johnson graph for some integers $n>k>0$ and a set $L=\{\ell_{1},\ell_{2},\ell_{3},\ldots,\ell_{s}\}\subset[0,k-1]$ with $\ell_{1}<\ell_{2}<\ldots<\ell_{s}$. Suppose $L$ contains $b$ full runs of consecutive integers. Then, as $n\rightarrow\infty$, there is a constant $c$ depending on $k$ and $L$ but not $n$ such that
@@
and a constant $C$ depending on $k,L$ and $L^{C}:=[0,k-1]\setminus L$ but not $n$ such that

### v3

@@ 2.2 Analyzing the linear program for the Lovász number of generalized Johnson graphs
$\sum_{\lambda\in\Lambda}\text{sgn}(\lambda)\prod_{i=m+1}^{m+p}P_{(i,\lambda(i))}=(-1)^{p}\frac{\prod_{i=1}^{p}(\ell_{m+i}+1)}{p!\prod_{i=1}^{p}(k-\ell_{m+i}-1)!}n^{pk-(\sum_{i=1}^{p}\ell_{m+i})-p}+O(n^{pk-(\sum_{i=1}^{p}\ell_{m+i})-p-1}).$
@@
We proceed by strong induction on $p$. The case $p=1$ follows from (5), as $\lambda(m+1)=m+1$, so $\text{sgn}(\lambda)=1$ and $P_{(m+1,\lambda(m+1))}=(-1)\frac{\ell_{m+1}+1}{(k-\ell_{m+1}-1)!}n^{k-\ell_{m+1}-1}+O(n^{k-\ell_{m+1}-2})$. Now suppose the statement of the lemma holds for all $11\leq r\leq p-$ 1; that is, assume that the lemma holds for any run of $r$ consecutive integers $\{\ell_{y},\ell_{y+1},\ldots,\ell_{y+r-1}\}$ in $L$. We show that the statement holds for $L_{m}$ under this assumption. Divide the permutations $\lambda$ in $\Lambda$ up according to the value of $j$ for which $\lambda(m+j)=m+1$. It follows from the conditions on $\lambda$ that $\lambda(m+i)=m+i+1$ for $1\leq i<j$, so that for these permutations $\lambda$,
$\prod_{i=m+1}^{m+j}P_{(i,\lambda(i))}=\binom{\ell_{m+1}+j}{\ell_{m+1}}(-1)^{\ell_{m+j}+1-\ell_{m+1}}\prod_{i=1}^{j}\frac{1}{(k-\ell_{m+i}-1)!}n^{jk-\sum_{i=1}^{j}\ell_{m+i}-j}+O(n^{jk-\sum_{i=1}^{j}\ell_{m+i}-j})$
$=(-1)^{j}\frac{\prod_{i=1}^{j}(\ell_{m+i}+1)}{j!\prod_{i=1}^{j}(k-\ell_{m+i}-1)!}n^{jk-\sum_{i=1}^{j}\ell_{m+i}-j}+O(n^{jk-\sum_{i=1}^{j}\ell_{m+i}-j}).$
@@
$\sum_{\begin{subarray}{c}\lambda\in\Lambda\\ \lambda(m+j)=m+1\end{subarray}}\text{sgn}(\lambda_{[m+j+1,m+p]})\prod_{i=m+j+1}^{m+p}P_{(i,\lambda(i))}$
$=(-1)^{p-j}\frac{1}{(p-j)!}\frac{\prod_{i=j+1}^{p}(\ell_{m+i}+1)}{\prod_{i=j+1}^{p}(k-\ell_{m+i}-1)!}n^{(p-j)k-(\sum_{i=j+1}^{p}\ell_{m+i})-(p-j)}+O(n^{(p-j)k-(\sum_{i=j+1}^{p}\ell_{m+i})-(p-j)-1}),$
@@
$\sum_{\begin{subarray}{c}\lambda\in\Lambda\\ \lambda(m+j)=m+1\end{subarray}}\text{sgn}(\lambda)\prod_{i=m+1}^{m+p}P_{(i,\lambda(i))}$
@@
$(-1)^{p+j-1}\frac{1}{j!(p-j)!}\frac{\prod_{i=1}^{p}(\ell_{m+i}+1)}{\prod_{i=1}^{p}(k-\ell_{m+i}-1)!}n^{pk-(\ell_{m+1}+\cdots+\ell_{m+p})-p}+O(n^{pk-(\ell_{m+1}+\cdots+\ell_{m+p})-p-1}).$
@@
$\sum_{\lambda\in\Lambda}\text{sgn}(\lambda)\prod_{i=m+1}^{m+p}P_{(i,\lambda(i))}=\sum_{j=1}^{p}\sum_{\begin{subarray}{c}\lambda\in\Lambda\\ \lambda(m+j)=m+1\end{subarray}}\text{sgn}(\lambda)\prod_{i=m+1}^{m+p}P_{(i,\lambda(i))}$
@@
$\sum_{\sigma\in\Sigma}\text{sgn}(\sigma)\prod_{i=1}^{s}P_{(i,\sigma(i))}=\prod_{i=1}^{b}\left(\sum_{\sigma_{r_{i}}}\text{sgn}(\sigma_{r_{i}})\prod_{a=c_{i}}^{c_{i}+m_{i}-1}P_{(a,\sigma_{r_{i}}(a))}\right)$
@@
For each $1\leq i\leq s$, let $a_{k-\ell_{i}}$ be the expression given by (4). Then, $a_{k-\ell_{i}}$ is a rational expression in $n$ of degree $s-i+1$ with positive leading order term $C_{i}$, so that
@@
${\bf v}=P^{-1}\begin{pmatrix}0\\ 0\\ \vdots\\ 0\\ -1\end{pmatrix}.$
Therefore, we only need to analyze the terms in column $s$ of $P^{-1}$. We first determine the leading order term of $P_{(1,s)}^{-1}$. We have $P_{(1,s)}^{-1}=\frac{(-1)^{s+1}}{\text{det}(P)}\text{det}(P^{(s,1)})$. We find the leading order term of $\text{det}(P^{(s,1)})$ by the Leibniz formula. Let $\sigma$ be a permutation which contributes to the leading order term of $\text{det}(P^{(s,1)})$. We claim that $\sigma$ must be the identity permutation, i.e., $\sigma(m)=m$ for $11\leq m\leq s-$ 1. By (5), the leading order term in the first row is $\frac{1}{(k-\ell_{2})!}\binom{k-\ell_{1}-1}{k-\ell_{2}}n^{k-\ell_{2}}$, as $P_{(1,i)}^{(s,1)}=\Theta(n^{k-\ell_{i+1}})$ for $11\leq i\leq s-$ 1, so in particular $\sigma(1)=1$. To show $\sigma(j)=j$ for $j>1$, observe that if $\sigma(m)=m$ for $1\leq m<j$, then $\sigma(j)\geq j$, and again by (5) we have $P_{(j,i)}^{(s,1)}=\Theta(n^{k-\ell_{i+1}})$ for $j\leq i\leq s-1$. Thus, $\sigma(j)=j$, and by induction we must have $\sigma(m)=m$ for $11\leq m\leq s-$ 1. Thus, by the Leibniz formula
$\text{det}(P^{(s,1)})$ $=\prod_{m=1}^{s-1}P_{(m,m)}^{(s,1)}+O(n^{(s-1)k-\sum_{i=2}^{s}\ell_{i}-1})$
$=\binom{k-\ell_{1}-1}{k-\ell_{2}}\cdots\binom{k-\ell_{s-1}-1}{k-\ell_{s}}\frac{1}{\prod_{i=2}^{s}(k-\ell_{i})!}n^{(s-1)k-\sum_{i=2}^{s}\ell_{i}}+O(n^{(s-1)k-\sum_{i=2}^{s}\ell_{i}-1}).$
@@
$a_{k-\ell_{1}}=\frac{(-1)\binom{k}{k-\ell_{1}}\binom{n-k}{k-\ell_{1}}(-1)^{s+1}\text{det}(P^{(s,1)})}{\text{det}(P)}$
@@
Now, let $2\leq i\leq s$. We have $P_{(i,s)}^{-1}=\frac{(-1)^{s+i}}{\text{det}(P)}\text{det}(P^{(s,i)})$, so we need to find the leading order term of $\text{det}(P^{(s,i)})$ by the Leibniz formula. Let $\Sigma^{*}$ denote the set of permutations of $[s-1]$ which contribute to the leading order term of the Leibniz formula for $\text{det}(P^{(s,i)})$. We claim that the permutations in $\Sigma^{*}$ permute the elements in $[1,i-1]$ and the elements in $[i,s-1]$ separately.
Claim 2.12.
Let $\sigma\in\Sigma^{*}$ be a permutation which contributes to the leading order term of $\text{det}(P^{(s,i)})$ in the Leibniz formula. Then, $\sigma(j)\leq i-1$ for $11\leq j\leq i-$ 1.
Proof of Claim 2.12.
Let $\chi$ be a permutation of $[s-1]$ which has $\chi(j)=z$ for some $11\leq j\leq i-$ 1 and some $z\geq i$. Then, there is some $m\geq i$ and some $11\leq b\leq i-$ 1 such that $\chi(m)=b$. Let $\tau$ be the permutation obtained from $\chi$ by swapping $b$ and $z$, so that $\tau(j)=b$, $\tau(m)=z$ and $\tau(p)=\chi(p)$ for $p\in[s-1]\setminus\{j,m\}$. We show that
$\prod_{w=1}^{s-1}P_{(w,\chi(w))}^{(s,i)}=o\left(\prod_{w=1}^{s-1}P_{(w,\tau(w))}^{(s,i)}\right),$
which implies that $\chi\notin\Sigma^{*}$.
Since $\tau$ and $\chi$ agree except on $j$ and $m$, it is enough to compare $P_{(j,\chi(j))}P_{(m,\chi(m))}$ and $P_{(j,\tau(j))}P_{(m,\tau(m))}$. By (5), we have that $P_{(j,z)}^{(s,i)}=\Theta(n^{k-\ell_{z+1}})$ and $P_{(m,b)}^{(s,i)}=\Theta(n^{k-\ell_{m}-1})$, so that
$P_{(j,\chi(j))}P_{(m,\chi(m))}=\Theta(n^{2k-\ell_{z+1}-\ell_{m}-1}).$
For the permutation $\tau$, there are four cases: if $j<b$, then $P_{(j,b)}^{(s,i)}=\Theta(n^{k-\ell_{b}})$; otherwise, $P_{(j,b)}^{(s,i)}=\Theta(n^{k-\ell_{j}-1})$. Similarly, if $m<z+1$, then $P_{(m,z)}^{(s,i)}=\Theta(n^{k-\ell_{z+1}})$, and otherwise $P_{(m,z)}^{(s,i)}=\Theta(n^{k-\ell_{m}-1})$. In each of the four cases, there is one term of order $\Theta(n^{k-\ell_{z+1}})$ or order $\Theta(n^{k-\ell_{m}-1})$ and one term of order $\Theta(n^{k-\ell_{b}})$ or $\Theta(n^{k-\ell_{j}-1})$. Since $\max\{j,b\}\leq i-1$ and $\min\{m,z\}\geq i$, it follows that $\min\{\ell_{m}+1,\ell_{z+1}\}>\max\{\ell_{b},\ell_{j}+1\}$, implying that
$P_{(j,\chi(j))}P_{(m,\chi(m))}=o(P_{(j,\tau(j))}P_{(m,\tau(m))}),$
completing the proof of the claim. ∎
For a permutation $\sigma\in\Sigma^{*}$, let $\sigma_{[1,i-1]}$ be the restriction of $\sigma$ to the set $[1,i-1]$ and let $\sigma_{[i,s-1]}$ be the restriction of $\sigma$ to the set $[i,s-1]$. Note that $|\text{inv}(\sigma)|=|\text{inv}(\sigma_{[1,i-1]})|+|\text{inv}(\sigma_{[i,s-1]})|$, so that
$\sum_{\sigma\in\Sigma^{*}}\text{sgn}(\sigma)\prod_{i=1}^{s-1}P_{(i,\sigma(i))}=\left(\sum_{\sigma_{[1,i-1]}}\text{sgn}(\sigma_{[1,i-1]})\prod_{m=1}^{i-1}P_{(m,\sigma(m))}^{(s,i)}\right)\cdot\left(\sum_{\sigma_{[i,s-1]}}\text{sgn}(\sigma_{[i,s-1]})\prod_{m=i}^{s-1}P_{(m,\sigma(m))}^{(s,i)}\right).$
We first consider the restriction of the permutation $\sigma$ to the set $[1,i-1]$. The $(i-1)\times{(i-1)}$ submatrix $P_{i-1}$ of $P^{(s,i)}$ given by the first $i-1$ rows and the first $i-1$ columns of $P^{(s,i)}$ is equivalent to the matrix obtained from the $L$-system with parameters $k$ and $L_{i-1}:=\{\ell_{1},\ldots,\ell_{i-1}\}$:
$P_{i-1}=\begin{pmatrix}P_{k-\ell_{1}}^{\ell_{1}+1}&P_{k-\ell_{2}}^{\ell_{1}+1}&\ldots&P_{k-\ell_{i-1}}^{\ell_{1}+1}\\ \vdots&\vdots&&\vdots\\ P_{k-\ell_{1}}^{\ell_{i-1}+1}&P_{k-\ell_{2}}^{\ell_{i-1}+1}&\ldots&P_{k-\ell_{i-1}}^{\ell_{i-1}+1}\end{pmatrix}.$
Hence, by the proof of Lemma 2.8, we have that
$\sum_{\sigma_{[1,i-1]}}\text{sgn}(\sigma_{[1,i-1]})\prod_{m=1}^{i-1}P_{(m,\sigma(m))}^{(s,i)}=(-1)^{i-1}D_{i}n^{k(i-1)-\sum_{j=1}^{i-1}\ell_{j}-(i-1)}+O(n^{k(i-1)-\sum_{j=1}^{i-1}\ell_{j}-(i-1)-1}),$
where
$D_{i}=\frac{\prod_{j=1}^{i-1}(\ell_{j}+1)}{\prod_{j=1}^{a}m_{j}^{\prime}!\prod_{j=1}^{i-1}(k-\ell_{j}-1)!},$
where for simplicity we have assumed $L_{i-1}$ has $a$ full runs of consecutive integers with lengths $m_{j}^{\prime}$ for $1\leq j\leq a$.
For the restriction of the permutation $\sigma$ to $[i,s-1]$, by a similar inductive argument as was used for $\text{det}(P^{(s,1)})$, the only permutation on $[i,s-1]$ which contributes to the leading order term in the Leibniz formula is the identity permutation on $[i,s-1]$, so
$\sum_{\sigma_{[i,s-1]}}\text{sgn}(\sigma_{[i,s-1]})\prod_{m=i}^{s-1}P_{(m,\sigma(m))}^{(s,i)}=\prod_{m=i}^{s-1}P_{(m,m)}^{(s,i)}=E_{i}n^{(s-i)k-(\ell_{i+1}+\ldots+\ell_{s})}+O(n^{(s-i)k-(\ell_{i+1}+\ldots+\ell_{s})-1}),$
where
$E_{i}=\binom{k-\ell_{i}-1}{k-\ell_{i+1}}\cdots\binom{k-\ell_{s-1}-1}{k-\ell_{s}}\frac{1}{\prod_{j=i+1}^{s}(k-\ell_{j})!}.$
These calculations give that
$\text{det}(P^{(s,i)})$ $=\sum_{\sigma\in\Sigma^{*}}\text{sgn}(\sigma)\prod_{m=1}^{s-1}P_{(m,\sigma(m))}^{(s,i)}+O(n^{(s-1)k-(\sum_{j=1,j\neq i}^{s}\ell_{j})-(i-1)-1})$
$=(-1)^{i-1}D_{i}E_{i}n^{(s-1)k-(\sum_{j=1,j\neq i}^{s}\ell_{j})-(i-1)}+O(n^{(s-1)k-(\sum_{j=1,j\neq i}^{s}\ell_{j})-(i-1)-1}).$
Hence, by Lemma 2.8,
$a_{k-\ell_{i}}=\frac{(-1)\binom{k}{k-\ell_{i}}\binom{n-k}{k-\ell_{i}}}{\text{det}(P)}(-1)^{s+i}\text{det}(P^{(s,i)})=C_{i}n^{s-i+1}-O(n^{s-i}),$
where
$C_{i}=\binom{k}{k-\ell_{i}}\frac{D_{i}E_{i}}{(k-\ell_{i})!C}$
is a positive constant (here $C$ is the positive constant in the statement of Lemma 2.8). ∎
@@
The inequalities (8) with $\ell_{m}+2\leq u\leq\ell_{m+1}$ for $11\leq m\leq s-$ 1 will be satisfied for large $n$ as the leading order term on the left-hand side is $a_{k-\ell_{m+1}}\frac{\binom{k-u}{k-\ell_{m}}}{\binom{k}{k-\ell_{m}}}=C_{m+1}\frac{\binom{k-u}{k-\ell_{m+1}}}{\binom{k}{k-\ell_{m+1}}}n^{s-m}-O(n^{s-m-1})$.
@@
It would be interesting to strengthen Theorem 1.2 by finding the exact value of $\vartheta(G(n,k,L))$ (as Theorem 1.5 does if $|L|=1$ or $|L|=k-1$) as $n\rightarrow\infty$. We have opted for the computations presented here, as they are sufficient to obtain the order of magnitude of $\vartheta(G(n,k,L))$ and the correct constant in front of the $n^{s}$ term.
We note that the argument used to prove Theorem 1.2 also asymptotically determines the best possible result the Delsarte linear programming bound [Del73, Sch79] can give for the graphs $G(n,k,L)$. Since the graphs $G(n,k,L)$ are unions of graphs in the Johnson scheme, Delsarte’s linear programming bound $\sigma(G(n,k,L))$ is equal to Schrijver’s variant of the Lovász number [Sch79], which is the same linear program as the linear program for the Lovász number in (1), with the additional restriction that the variables $a_{0},a_{1},\ldots,a_{n}$ must all be nonnegative. The solution in the proof of Theorem 1.2 has all $a_{k-\ell_{i}}$ nonnegative, so (4) is also a feasible solution for Delsarte’s linear programming bound, so as $n\rightarrow\infty$,
@@
Let $G=G(n,k,L)$ be a generalized Johnson graph for some integers $n>k>0$ and a set $L=\{\ell_{1},\ell_{2},\ell_{3},\ldots,\ell_{s}\}\subset[0,k-1]$ with $\ell_{1}<\ell_{2}<\ldots<\ell_{s}$. Suppose $L$ contains $b$ full runs of consecutive integers. Then, there is a constant $c$ depending on $k$ and $L$, but not on $n$ such that
@@
and a constant $C$ depending on $k$ and $L$ but not on $n$ such that